OpenStudy (anonymous):

-2.5x^2+2825x-115812.5=500000

OpenStudy (anonymous):

do you have a graphing calculator?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

x = 835 0r x = 295

OpenStudy (saifoo.khan):

here u go! (:

OpenStudy (anonymous):

you can easily find answers to any questions of this kind simply by graphing the entire equation in your calculator. When the graph appears, simply use the function that finds where the graph crosses the x-axis(x=0) and thats your answer. there can be multiple answers

OpenStudy (anonymous):

-2.5x^2+2825x-115812.5=500000 Rationalize the decimal fraction coefficients.\[-\frac{5}{2}\left(x^2\right)+2825x-\frac{231625}{2}=500000 \]Multiply both equation sides by -2/5, move the RHS to the LHS and simplify:\[x^2-1130 x+246325=0 \]The equation can be factored.\[(x-835) (x-295)=0 \]@saifoo.khan listed the solutions in the postings above.